OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

m1m2=-1 m1=-3 from the equation so m2=1/3 so perpendicular line's equation on (6,-1) is y+1=1/3(x-6) x-3y=9 ~ANS